# Surveying the MLLM Landscape: A Meta-Review of Current Surveys

### v1

@@ 2 Classical Natural Language Processing Methods
Early language model research was primarily based on the Markov process. Markov chains were used to describe random processes and sequence data, introducing basic methods for handling sequential data[1, 2]. Subsequently, Hidden Markov Models (HMMs) were proposed to perform continuous speech recognition via maximum likelihood estimation, dealing with hidden states in sequential data[3]. In the 1990s, n-gram-based statistical language models were widely used. N-gram models predicted the next word based on the context, establishing a simple yet effective foundation for language models[4]. With increased computational power, neural networks began to be applied in NLP. In 1997, Long Short-Term Memory (LSTM) networks were proposed to solve the problem of long-term dependencies using memory cells, making RNNs more effective in handling long sequential data[5, 6]. In 1998, Yann LeCun and others further developed RNNs for document recognition, capable of handling sequential data[7]. Word embedding techniques achieved significant breakthroughs in the early 2010s. In 2003, the Word2Vec model was introduced, capturing semantic relationships between words through efficient word vector training methods[8]. This method introduced the Skip-gram and CBOW models, greatly improving the quality of word vectors.
@@
In 2009, ImageNet, a large visual database containing over 14 million annotated images, was launched for object recognition and image classification tasks[9]. The advent of ImageNet provided deep learning with abundant data resources, making it possible to train deep neural networks on large-scale datasets. In 2012, AlexNet’s success in the ImageNet[9] competition marked the resurgence of deep learning[10]. AlexNet significantly improved image classification performance through deep convolutional neural networks, bringing deep learning back to the forefront of research and inspiring researchers to explore its potential in other fields like NLP. In 2014, the GloVe model trained word vectors using global word co-occurrence matrices, further improving semantic representation of words[11]. Additionally, the sequence-to-sequence learning (Seq2Seq) model was proposed for machine translation, handling sequence transformation tasks through an encoder-decoder structure[12]. In 2015, Bahdanau et al. introduced the attention mechanism, which dynamically aligned source and target sequences, improving machine translation performance[13]. In 2016, the Byte Pair Encoding (BPE) algorithm was proposed to handle rare and out-of-vocabulary words through subword units, significantly enhancing model generalization ability[14]. In 2018, the ELMo (Embeddings from Language Models) model was introduced, capturing contextual information through bidirectional LSTM (Long Short-Term Memory) networks. Specifically, ELMo used two independently trained LSTMs, one processing text from left to right and the other from right to left, concatenating the hidden states of these two LSTMs to form the final word representation. This method improved the quality of word representations by capturing lexical semantics and syntactic information while fully utilizing contextual information[15].
@@ 3.2 Transformer
Vaswani et al. introduced the Transformer, a deep learning model widely used for natural language processing (NLP) tasks [16]. As the cornerstone of large-scale language models (LLMs) today, the core of the Transformer model is the attention mechanism, which captures relationships between different positions in the input sequence without relying on traditional recurrent neural network (RNN) structures. This enables parallel computation and improves training efficiency.
@@ 3.2.1 BERT and GPT
BERT (Bidirectional Encoder Representations from Transformers) is a pre-training model proposed by Devlin et al.[17]. The core idea of BERT is to generate contextual representations of words using a bidirectional Transformer encoder. It achieves performance improvements across various NLP tasks by performing unsupervised pre-training on large-scale corpora, followed by fine-tuning on specific tasks. BERT’s pre-training involves two tasks: Masked Language Model (MLM) and Next Sentence Prediction (NSP).
@@
GPT (Generative Pre-trained Transformer) is a series of generative pre-training models proposed by OpenAI[18, 19, 20]. Unlike BERT, GPT mainly uses a unidirectional Transformer decoder to generate text. GPT’s training consists of two stages: first, pre-training on large-scale unlabelled data, and then fine-tuning on specific tasks. During the pre-training stage, the model uses an autoregressive method to predict the next word, thereby learning the sequence information of words.
@@ 3.3 Multimodal Large Language Models (MLLM)
In the context of aligning text and images, the initial challenge was the necessity of vast amounts of data to ensure models could effectively understand and correlate different modalities. Early methods relied on manually labeled data, which was both costly and inefficient. Some of these early methods include DeViSE (A Deep Visual-Semantic Embedding Model) [21], VSE (Visual Semantic Embeddings) [22] and CCA (Canonical Correlation Analysis) [23]. These models attempted to align images and texts in a shared representation space using labeled datasets, but their performance was limited due to the restricted size of the data.
@@ 3.3.1 Contrastive Learning-based Multimodal Alignment
With the advent of contrastive learning, this field has significant progress. The core idea of contrastive learning is to bring similar sample pairs (such as related image and text pairs) closer together while pushing dissimilar pairs apart in the shared representation space, thus achieving multimodal alignment. Early contrastive learning methods focused primarily on the visual domain, such as MoCo (Momentum Contrast) [24, 25], SimCLR (Simple Framework for Contrastive Learning of Visual Representations) [26, 27], Instance Discrimination [28], CPC (Contrastive Predictive Coding) [29], CMC (Contrastive Multiview Coding) [30], PIRL (Pretext-Invariant Representation Learning) [31], SwAV (Swapping Assignments between Multiple Views of the Same Image) [32], and BYOL (Bootstrap Your Own Latent) [33]. These methods leveraged large unlabeled image datasets and used contrastive loss functions to significantly enhance image feature representations.
In the realm of multimodal data alignment, contrastive learning methods have made remarkable strides. CLIP (Contrastive Language-Image Pre-Training) [34] represents a major breakthrough in this area. CLIP leverages a large-scale dataset of image-text pairs and employs a contrastive loss function to achieve robust text-image alignment. Specifically, CLIP encodes images and texts into a shared representation space to align them. Given an image $I$ and its corresponding text description $T$, the image encoder $f(\cdot)$ and text encoder $g(\cdot)$ map them into a common representation space: the image feature vector $f(I)$ and the text feature vector $g(T)$. In this space, matched image-text pairs exhibit high similarity, while mismatched pairs show low similarity. Pretraining on hundreds of millions of image-text pairs allows CLIP to learn rich cross-modal representations, excelling in downstream tasks.
@@
There are numerous models that employ CLIP visual or text encoders, demonstrating their broad applicability in multimodal alignment. For instance, METER (Multimodal End-to-end TransformER) [35] integrates CLIP encoders to align image and text representations and employs optimal transport techniques to improve the performance of vision-language tasks. SimVLM (Simple Visual Language Model) [36] uses CLIP encoders for weakly supervised pretraining, achieving excellent performance in vision-language tasks. BLIP-2 (Bootstrapping Language-Image Pre-training) [37] utilizes frozen CLIP image encoders with large language models to improve image-text pretraining. The Stable Diffusion series, including high-resolution image synthesis [38], personalized stylization [39], and improved high-resolution image synthesis [40], also employs CLIP encoders to optimize image generation and super-resolution tasks.
@@ 3.3.2 Model-Based Techniques for Multimodal Alignment
Model-based techniques for multimodal alignment aim to create unified representations of different modalities, such as images and text, by leveraging large-scale models that handle both vision and language tasks. ALIGN (Large-scale Image-Language Pretraining) [41] employs a dual-encoder architecture that aligns images and text in a shared embedding space using contrastive learning, significantly improving zero-shot image classification and retrieval. Similarly, Florence [42] incorporates a vision transformer model with a text encoder to align image and text modalities, demonstrating superior performance in visual understanding tasks. These models focus on pretraining using vast amounts of weakly labeled image-text data, contributing to better generalization across a range of vision-language benchmarks.
Other models utilize fine-grained alignment strategies to enhance performance in more specialized tasks. UNITER (UNiversal Image-TExt Representation) [43] jointly learns image-text alignments using a transformer-based architecture, introducing a unique masked language modeling task to improve contextual understanding between modalities. On the other hand, ViLT (Vision-and-Language Transformer) [44] reduces the complexity of multimodal fusion by directly feeding visual patch embeddings into a transformer, bypassing the need for separate image encoders, while still maintaining competitive performance in tasks like image-text retrieval and visual question answering. These model-based techniques highlight the versatility and scalability of transformer-based approaches in multimodal alignment tasks, addressing both high-level semantic matching and fine-grained object-level alignment.
{forest}
[…]
Shi et al.  2024, tnode] ] [Data, xnode, l sep=7mm [Bai et al.  2024
Yang et al.  2024, tnode] ] [Efficient, xnode, l sep=7mm [Xu et al.  2024
@@
Li et al.  2023, tnode] ] [RAG/info retrieval, xnode, l sep=7mm [Author Y et al.  2022
@@
Chen et al.  2024, tnode] ] [Evaluation, xnode, l sep=7mm [Lu et al.  2024
@@
Chen et al.  2024, tnode] ] [Bias, xnode, l sep=7mm [Liang et al.  2024
@@
Li et al.  2023, tnode] ] [Agent, xnode, l sep=7mm [Wang et al.  2024
@@
Xu et al.  2024, tnode] ] [Security, xnode, l sep=7mm [Caffagni et al.  2024
@@
Pi et al.  2024, tnode] ] [Application, xnode, l sep=7mm [Minaee et al.  2024
@@
Zhang et al.  2023, tnode] ] [General, xnode, l sep=7mm [Caffagni et al.  2024
@@ 4.2 Applications and Agents
The survey [45] focuses on the integration of LLMs into the legal domain, where they have been used for tasks such as legal advice, document processing, and judicial assistance. The key trend here is the application of specialized LLMs, such as LawGPT and LexiLaw, which address the nuances of legal language and reasoning. One significant challenge identified is maintaining judicial independence and addressing the ethical implications of biased data in legal decision-making.
@@
In the field of autonomous driving, MLLMs are increasingly being used to improve perception, decision-making, and human-vehicle interaction, as noted in [46]. A key trend is the integration of multimodal data, such as LiDAR, maps, and images, which enhance the vehicle’s ability to process complex driving environments. However, challenges related to real-time data processing and ensuring safety in diverse driving conditions remain significant.
@@
[47] explores the application of LLMs in solving mathematical tasks, such as calculation and reasoning. Techniques like Chain-of-Thought (CoT) prompting have significantly improved model performance in complex mathematical problems. However, the scarcity of high-quality datasets and the complexity of mathematical reasoning pose ongoing challenges.
@@
The healthcare survey [48] reviews the use of multimodal learning for tasks like image fusion, report generation, and cross-modal retrieval. A key trend is the rise of foundation models such as GPT-4 and CLIP in processing medical data. Despite advancements, these models have not yet achieved universal intelligence, and concerns related to data integration and ethical considerations remain major barriers.
@@
The use of LLMs in robotics [49] focuses on their ability to improve perception, decision-making, and control in robotic systems. The main trend identified is the potential for LLMs to advance embodied intelligence, where robots understand and interact with the physical world. However, challenges related to real-time perception, control, and integration with existing robotic technologies persist.
@@
In multilingual settings, the ability of LLMs to process multiple languages is highlighted in [50]. There is significant progress in handling multiple languages, but low-resource languages and security issues in multilingual models remain challenges. Emerging techniques like multilingual Chain-of-Thought reasoning show promise for future development.
@@
The survey [51] focuses on the role of LLMs in gaming, particularly in generating dynamic Non-Player Characters (NPCs), enhancing player interaction, and even assisting in game design. The challenge of hallucinations, where LLMs generate plausible but incorrect outputs, is a major limitation in real-time gaming environments. Improving context management and memory within gaming systems is a future research priority.
@@
[52] discusses the application of LLMs in audio processing tasks, such as Automatic Speech Recognition (ASR) and music generation. The integration of multimodal data from speech, music, and environmental sounds into a single model, like SeamlessM4T, marks a significant trend. However, scalability and data diversity remain issues to address.
@@
In [53], the focus is on video understanding through the use of Vid-LLMs (Video-Large Language Models), which combine video, text, and audio inputs to analyze and understand video content. While these models are promising for tasks like video summarization and captioning, challenges related to processing long videos and maintaining contextual coherence need further exploration.
@@
The use of LLMs for citation tasks is discussed in [54], where LLMs significantly improve citation recommendation, classification, and summarization tasks. Additionally, citation data enhances LLM performance by enabling multi-hop knowledge across documents. Future research needs to address the expansion of citation networks and integration of non-academic sources.

### v2

@@ 2 Classical Natural Language Processing Methods
Early language model research was primarily based on the Markov process. Markov chains were used to describe random processes and sequence data, introducing basic methods for handling sequential data[dynkin1965markov, chung1967markov]. Subsequently, Hidden Markov Models (HMMs) were proposed to perform continuous speech recognition via maximum likelihood estimation, dealing with hidden states in sequential data[bahl1983maximum]. In the 1990s, n-gram-based statistical language models were widely used. N-gram models predicted the next word based on the context, establishing a simple yet effective foundation for language models[jurafsky2000speech]. With increased computational power, neural networks began to be applied in NLP. In 1997, Long Short-Term Memory (LSTM) networks were proposed to solve the problem of long-term dependencies using memory cells, making RNNs more effective in handling long sequential data[grossberg2013recurrent, hochreiter1997long]. In 1998, Yann LeCun and others further developed RNNs for document recognition, capable of handling sequential data[lecun1998gradient]. Word embedding techniques achieved significant breakthroughs in the early 2010s. In 2003, the Word2Vec model was introduced, capturing semantic relationships between words through efficient word vector training methods[mikolov2013efficient]. This method introduced the Skip-gram and CBOW models, greatly improving the quality of word vectors.
@@
In 2009, ImageNet, a large visual database containing over 14 million annotated images, was launched for object recognition and image classification tasks[deng2009imagenet]. The advent of ImageNet provided deep learning with abundant data resources, making it possible to train deep neural networks on large-scale datasets. In 2012, AlexNet’s success in the ImageNet[deng2009imagenet] competition marked the resurgence of deep learning[krizhevsky2012imagenet]. AlexNet significantly improved image classification performance through deep convolutional neural networks, bringing deep learning back to the forefront of research and inspiring researchers to explore its potential in other fields like NLP. In 2014, the GloVe model trained word vectors using global word co-occurrence matrices, further improving semantic representation of words[pennington2014glove]. Additionally, the sequence-to-sequence learning (Seq2Seq) model was proposed for machine translation, handling sequence transformation tasks through an encoder-decoder structure[sutskever2014sequence]. In 2015, Bahdanau et al. introduced the attention mechanism, which dynamically aligned source and target sequences, improving machine translation performance[bahdanau2014neural]. In 2016, the Byte Pair Encoding (BPE) algorithm was proposed to handle rare and out-of-vocabulary words through subword units, significantly enhancing model generalization ability[sennrich2015neural]. In 2018, the ELMo (Embeddings from Language Models) model was introduced, capturing contextual information through bidirectional LSTM (Long Short-Term Memory) networks. Specifically, ELMo used two independently trained LSTMs, one processing text from left to right and the other from right to left, concatenating the hidden states of these two LSTMs to form the final word representation. This method improved the quality of word representations by capturing lexical semantics and syntactic information while fully utilizing contextual information[peters2018deep].
@@ 3.2 Transformer
Vaswani et al. introduced the Transformer, a deep learning model widely used for natural language processing (NLP) tasks [vaswani2017attention]. As the cornerstone of large-scale language models (LLMs) today, the core of the Transformer model is the attention mechanism, which captures relationships between different positions in the input sequence without relying on traditional recurrent neural network (RNN) structures. This enables parallel computation and improves training efficiency.
@@ 3.2.1 BERT and GPT
BERT (Bidirectional Encoder Representations from Transformers) is a pre-training model proposed by Devlin et al.[devlin2018bert]. The core idea of BERT is to generate contextual representations of words using a bidirectional Transformer encoder. It achieves performance improvements across various NLP tasks by performing unsupervised pre-training on large-scale corpora, followed by fine-tuning on specific tasks. BERT’s pre-training involves two tasks: Masked Language Model (MLM) and Next Sentence Prediction (NSP).
@@
GPT (Generative Pre-trained Transformer) is a series of generative pre-training models proposed by OpenAI[radford2018improving, radford2019language, brown2020language]. Unlike BERT, GPT mainly uses a unidirectional Transformer decoder to generate text. GPT’s training consists of two stages: first, pre-training on large-scale unlabelled data, and then fine-tuning on specific tasks. During the pre-training stage, the model uses an autoregressive method to predict the next word, thereby learning the sequence information of words.
@@ 3.3 Multimodal Large Language Models (MLLM)
In the context of aligning text and images, the initial challenge was the necessity of vast amounts of data to ensure models could effectively understand and correlate different modalities. Early methods relied on manually labeled data, which was both costly and inefficient. Some of these early methods include DeViSE (A Deep Visual-Semantic Embedding Model) [frome2013devise], VSE (Visual Semantic Embeddings) [kiros2014unifying] and CCA (Canonical Correlation Analysis) [hardoon2004canonical]. These models attempted to align images and texts in a shared representation space using labeled datasets, but their performance was limited due to the restricted size of the data.
@@ 3.3.1 Contrastive Learning-based Multimodal Alignment
With the advent of contrastive learning, this field has significant progress. The core idea of contrastive learning is to bring similar sample pairs (such as related image and text pairs) closer together while pushing dissimilar pairs apart in the shared representation space, thus achieving multimodal alignment. Early contrastive learning methods focused primarily on the visual domain, such as MoCo (Momentum Contrast) [he2020momentum, chen2020improved], SimCLR (Simple Framework for Contrastive Learning of Visual Representations) [chen2020simple, chen2020big], Instance Discrimination [wu2018unsupervised], CPC (Contrastive Predictive Coding) [oord2018representation], CMC (Contrastive Multiview Coding) [tian2020contrastive], PIRL (Pretext-Invariant Representation Learning) [misra2020self], SwAV (Swapping Assignments between Multiple Views of the Same Image) [caron2020unsupervised], and BYOL (Bootstrap Your Own Latent) [grill2020bootstrap]. These methods leveraged large unlabeled image datasets and used contrastive loss functions to significantly enhance image feature representations.
In the realm of multimodal data alignment, contrastive learning methods have made remarkable strides. CLIP (Contrastive Language-Image Pre-Training) [radford2021learning] represents a major breakthrough in this area. CLIP leverages a large-scale dataset of image-text pairs and employs a contrastive loss function to achieve robust text-image alignment. Specifically, CLIP encodes images and texts into a shared representation space to align them. Given an image $I$ and its corresponding text description $T$, the image encoder $f(\cdot)$ and text encoder $g(\cdot)$ map them into a common representation space: the image feature vector $f(I)$ and the text feature vector $g(T)$. In this space, matched image-text pairs exhibit high similarity, while mismatched pairs show low similarity. Pretraining on hundreds of millions of image-text pairs allows CLIP to learn rich cross-modal representations, excelling in downstream tasks.
@@
There are numerous models that employ CLIP visual or text encoders, demonstrating their broad applicability in multimodal alignment. For instance, METER (Multimodal End-to-end TransformER) [dou2022empirical] integrates CLIP encoders to align image and text representations and employs optimal transport techniques to improve the performance of vision-language tasks. SimVLM (Simple Visual Language Model) [wang2021simvlm] uses CLIP encoders for weakly supervised pretraining, achieving excellent performance in vision-language tasks. BLIP-2 (Bootstrapping Language-Image Pre-training) [li2023blip] utilizes frozen CLIP image encoders with large language models to improve image-text pretraining. The Stable Diffusion series, including high-resolution image synthesis [rombach2022high], personalized stylization [yang2023pixel], and improved high-resolution image synthesis [podell2023sdxl], also employs CLIP encoders to optimize image generation and super-resolution tasks.
@@ 3.3.2 Model-Based Techniques for Multimodal Alignment
Model-based techniques for multimodal alignment aim to create unified representations of different modalities, such as images and text, by leveraging large-scale models that handle both vision and language tasks. ALIGN (Large-scale Image-Language Pretraining) [DBLP:journals/corr/abs-2102-05918] employs a dual-encoder architecture that aligns images and text in a shared embedding space using contrastive learning, significantly improving zero-shot image classification and retrieval. Similarly, Florence [DBLP:journals/corr/abs-2111-11432] incorporates a vision transformer model with a text encoder to align image and text modalities, demonstrating superior performance in visual understanding tasks. These models focus on pretraining using vast amounts of weakly labeled image-text data, contributing to better generalization across a range of vision-language benchmarks.
Other models utilize fine-grained alignment strategies to enhance performance in more specialized tasks. UNITER (UNiversal Image-TExt Representation) [DBLP:journals/corr/abs-1909-11740] jointly learns image-text alignments using a transformer-based architecture, introducing a unique masked language modeling task to improve contextual understanding between modalities. On the other hand, ViLT (Vision-and-Language Transformer) [kim2021viltvisionandlanguagetransformerconvolution] reduces the complexity of multimodal fusion by directly feeding visual patch embeddings into a transformer, bypassing the need for separate image encoders, while still maintaining competitive performance in tasks like image-text retrieval and visual question answering. These model-based techniques highlight the versatility and scalability of transformer-based approaches in multimodal alignment tasks, addressing both high-level semantic matching and fine-grained object-level alignment.
Multimodal Large Language Models Continual Learning Feng et al.  2023
Shi et al.  2024
Data Bai et al.  2024
Yang et al.  2024
Efficient Xu et al.  2024
@@
Li et al.  2023
RAG/info retrieval Author Y et al.  2022
@@
Chen et al.  2024
Evaluation Lu et al.  2024
@@
Chen et al.  2024
Bias Liang et al.  2024
@@
Li et al.  2023
Agent Wang et al.  2024
@@
Xu et al.  2024
Security Caffagni et al.  2024
@@
Pi et al.  2024
Application Minaee et al.  2024
@@
Zhang et al.  2023
General Caffagni et al.  2024
@@ 4.2 Applications and Agents
The survey [minaee2024large] focuses on the integration of LLMs into the legal domain, where they have been used for tasks such as legal advice, document processing, and judicial assistance. The key trend here is the application of specialized LLMs, such as LawGPT and LexiLaw, which address the nuances of legal language and reasoning. One significant challenge identified is maintaining judicial independence and addressing the ethical implications of biased data in legal decision-making.
@@
In the field of autonomous driving, MLLMs are increasingly being used to improve perception, decision-making, and human-vehicle interaction, as noted in [cui2023surveymultimodallargelanguage]. A key trend is the integration of multimodal data, such as LiDAR, maps, and images, which enhance the vehicle’s ability to process complex driving environments. However, challenges related to real-time data processing and ensuring safety in diverse driving conditions remain significant.
@@
[liu2023mathematical] explores the application of LLMs in solving mathematical tasks, such as calculation and reasoning. Techniques like Chain-of-Thought (CoT) prompting have significantly improved model performance in complex mathematical problems. However, the scarcity of high-quality datasets and the complexity of mathematical reasoning pose ongoing challenges.
@@
The healthcare survey [lin2024multimodallearningdelivereduniversal] reviews the use of multimodal learning for tasks like image fusion, report generation, and cross-modal retrieval. A key trend is the rise of foundation models such as GPT-4 and CLIP in processing medical data. Despite advancements, these models have not yet achieved universal intelligence, and concerns related to data integration and ethical considerations remain major barriers.
@@
The use of LLMs in robotics [zeng2023large] focuses on their ability to improve perception, decision-making, and control in robotic systems. The main trend identified is the potential for LLMs to advance embodied intelligence, where robots understand and interact with the physical world. However, challenges related to real-time perception, control, and integration with existing robotic technologies persist.
@@
In multilingual settings, the ability of LLMs to process multiple languages is highlighted in [huang2024surveylargelanguagemodels]. There is significant progress in handling multiple languages, but low-resource languages and security issues in multilingual models remain challenges. Emerging techniques like multilingual Chain-of-Thought reasoning show promise for future development.
@@
The survey [gallotta2024large] focuses on the role of LLMs in gaming, particularly in generating dynamic Non-Player Characters (NPCs), enhancing player interaction, and even assisting in game design. The challenge of hallucinations, where LLMs generate plausible but incorrect outputs, is a major limitation in real-time gaming environments. Improving context management and memory within gaming systems is a future research priority.
@@
[latif2023sparks] discusses the application of LLMs in audio processing tasks, such as Automatic Speech Recognition (ASR) and music generation. The integration of multimodal data from speech, music, and environmental sounds into a single model, like SeamlessM4T, marks a significant trend. However, scalability and data diversity remain issues to address.
@@
In [tang2023video], the focus is on video understanding through the use of Vid-LLMs (Video-Large Language Models), which combine video, text, and audio inputs to analyze and understand video content. While these models are promising for tasks like video summarization and captioning, challenges related to processing long videos and maintaining contextual coherence need further exploration.
@@
The use of LLMs for citation tasks is discussed in [zhang2023large], where LLMs significantly improve citation recommendation, classification, and summarization tasks. Additionally, citation data enhances LLM performance by enabling multi-hop knowledge across documents. Future research needs to address the expansion of citation networks and integration of non-academic sources.